# Elucidating Mechanisms of Demographic Bias in LLMs for Healthcare

### v1

@@ Abstract
We know from prior work that LLMs encode social biases, and that this manifests in clinical tasks. In this work we adopt tools from mechanistic interpretability to unveil sociodemographic representations and biases within LLMs in the context of healthcare. Specifically, we ask: Can we identify activations within LLMs that encode sociodemographic information (e.g., gender, race)? We find that gender information is highly localized in middle MLP layers and can be reliably manipulated at inference time via patching. Such interventions can surgically alter generated clinical vignettes for specific conditions, and also influence downstream clinical predictions which correlate with gender, e.g., patient risk of depression. We find that representation of patient race is somewhat more distributed, but can also be intervened upon, to a degree. To our knowledge, this is the first application of mechanistic interpretability methods to LLMs for healthcare.
@@ 1 Introduction
In this work, we ask: Is the internal LLM encoding of patient demographics like gender and race localized? And, can we intervene upon this? To answer this we perform activation patching (Heimersheim and Nanda, 2024) on OLMo-7B-Instruct Groeneveld et al. (2024) in the context of clinical vignette generation. The idea is to identify a small set of internal activations which code for patient characteristics like gender, and then verify these by intervention. We offer the following contributions.
(1) We find that gender [^1] Following prior work Zack et al. (2024), we use “gender” instead of “sex” because of the ambiguity in how LLMs use the terms “male” and “female” (biological vs sociocultural). Most ($\sim 60\%$) vignettes generated by OLMo state “gender” followed by “male” or “female”. Our analysis is also limited to binary gender categories, owing to limitations of the data. information is localized to middle MLP layers; patching MLP activations of a single layer consistently alters patient gender in generated texts. Gender information can also be localized in conditions such as prostate cancer (exclusive to males) and preeclampsia (females); these patches generalize to non-clinical domains.
@@ 2.1 Vignette Generation
Is the encoding of patient gender localized or distributed in the LLM? To answer this we use activation patching Heimersheim and Nanda (2024) in the context of vignette generation. This requires access to model internals, so we adopt the open-source OLMo-7B-Instruct model in place of the opaque GPT-4. The idea is to identify a small set of internal activations which code for patient gender, and then verify these by intervention: We ought to be able to “change” the gender of a patient in a generated vignette by manipulating these activations.
@@
To find such activations, we first prompt the LLM to generate a vignette for a condition (strongly) associated with females (or males). Specifically, following prior work Zack et al. (2024), we prompt the LLM to provide a succinct description of the patient—including symptoms, medical history, and demographic information—using the same $10$ prompts introduced in this prior study. One of the prompts, e.g., is:
@@ 2.2 Activation Patching
[…]
Figure 2: Rewrite score distribution averaged over six conditions for which OLMo over-represents females (starting from $\ell$=$1$). Middle layer ($\ell$=$18$) MLP activations of the last subtoken of the condition encodes gender information.
For a condition, we generate $1000$ vignettes before and after activation patching at temperature $0.7$. We also experiment with scaling up the MLP activations by a factor, $c$. Figure 3(a) shows the ratio of male vignettes after patching for sarcoidosis and multiple sclerosis. Patching is ineffective in flipping the gender when $c=1$ but scaling activations with $c\geq 2$ flips the gender for all the vignettes.
@@
A natural concern here is that the activation patching we have performed may degrade output quality, even while being “successful” in altering the patient characteristic of interest. To assess if this is the case, we compute the perplexity using Llama-3.1-8B [^4] We use Llama as an external judge of perplexity rather than OLMo since the latter generated the text and so likely finds it high likelihood. Grattafiori et al. (2024) of $500$ vignettes LLM generated before and after patching. Figure 3(a) reports average perplexities as a function of $c$ for sarcoidosis and MS. Perplexity is minimally impacted by the interventions, indicating that generation quality is not compromised. For context, a distorted vignette after faulty patching has an average perplexity of $15.54$ (Appendix A.2).
@@ 2.3 Sexed Conditions
We have shown that we can extract a “male activation” which can consistently induce “maleness” via patching in clinical vignettes. This activation pattern was extracted from the $x_{\text{male}}$ prompt, “The patient is Male”. Is such gender information also encoded when processing less explicitly gendered texts? Here we consider the case of conditions which are inherently sexed. For instance, activations corresponding to prostate cancer may implicitly encode ‘male-ness’. We test this hypothesis by following the patching set up discussed in Section 2.2, but change $x_{\text{male}}$ from ‘The patient is Male’ to $x_{\text{vignette}}$ for ‘prostate cancer’. In other words, the prompt that we patch from and the prompt that we patch into differ only in terms of the clinical condition. We patch MLP activations of ‘prostate’ to the last sub-token of the condition in $x_{\text{vignette}}$ at layer $18$. We observe the same phenomenon as shown in Figure 3(a): patching after scaling up activations alters the gender $100\%$ of the time (Appendix A.3).
Patching to other domains
prompt pre-patching post-patching
The secretary slept because she had worked… he had worked…
@@
Figure 4: Rewrite score distribution for hepatitis B and sarcoidosis (starting from $\ell$=$1$). Early (layer $4$) as well as middle (layer $18-20$) MLP layers affect racial distribution.
@@ 3 Localizing Race
Figures 4(a) and 4(b) depict the average rewrite scores for hepatitis B and sarcoidosis respectively. Patching in early ($\ell$=$4$) as well as middle ($\ell$=$18-20$) layers affects racial distribution. Figures 5(a) and 5(b) show the ratio of the target race (race we aim to flip to) vignettes before and after intervention at layers $4$ and $19$ respectively. Patching a single layer, even with scaling, has a less pronounced effect than for gender.
Does patching multiple layers at a time help? We explore sliding window patching Meng et al. (2022), in which neighboring layers are simultaneously patched. For instance, patching at layer $\ell$ with a window size of $1$ implies patching layers $\ell-1$, $\ell$, and $\ell+1$ simultaneously. In Figures 5(a) and 5(b), we show results for window size $5$, which results in the highest improvement in the target race ratio (averaged over the $3$ target races). We report results of window sizes $1$ and $3$ in Tables 11 and 12. Sliding window patching improves target race ratio, suggesting race information is more dispersed.
Is generation quality impacted?
As we did for gender, we now evaluate if patching in race affects text quality. Specifically, we compute average perplexity under Llama-3.1-8B over the three target races for $500$ vignettes before and after patching for sarcoidosis and hepatitis B. Perplexities before and after patching are comparable (Figure 7).
@@ 4.1 Depression Risk
Similar to the setup in Sections 2 and 3, we extract the ‘male’ and ‘female’ representations using two simple source prompts, “The patient is Male” and “The patient is Female”, respectively. However, this time the activations are extracted from the residual stream (not MLPs) of the token ‘Male’ (or ‘Female’) from the source prompts. Also, the extracted activations are patched at the last token of the target prompt. See Appendix B for examples of the prompt. Our choice of last token is informed by and aligns with the claims of previous research Marks and Tegmark (2023); Todd et al. (2023) suggesting that the last token representation encodes information about the entire prompt. We patch the target prompt at layer $18$, and scale the activations with a factor of $2$; We picked these values using a validation set of $100$ BHCs.
@@ Results
Instruction-tuned LLMs are sensitive to instruction phrasings Sun et al. (2023); Ceballos-Arroyo et al. (2024). To ensure our results are robust, we perform the intervention on four different target prompts (see Appendix 10) to elicit risk of depression prediction. Table 5 reports the difference in risk prediction averaged over four prompts for each demographic. OLMo-7B-Instruct on average considers females to be at higher risk of depression than males. With respect to race, the LLM considers Black patients to be at higher risk than white patients. Table 4 shows some sample outputs. While implicit and explicit biases may manifest differently, we observe the same trend in disparity with explicit prompts as well (Appendix B).
@@ 4.2 Differential Diagnosis
We select one case each to study disparities between male/female patients and Black/white patients (see Appendix C for prompts and cases). We follow the set up described in Section 4.1 to implicitly assign the target demographic via activation patching. We use the same source prompts, target layers, and scaling factor.
@@ Results
We sample $500$ differential lists at temperature $0.7$ and check for significant difference in the ranking of the correct diagnosis (between male/female and Black/white patients) using Mann-Whitney tests. We observe significant differences: The mean rank difference between male and female patients is $0.24$ ($p=$$0.004$), and between Black and white patients this is $0.08$ ($p$=$0.02$). Figure 8 shows the rank distribution of the correct diagnosis when gender is explicitly stated and when it is patched; the trend in rank difference is similar whether gender is explicitly or implicitly assigned. (See Figure 11 in the Appendix for race).
@@ 5 Related Work
Recent works have shown that LLMs exhibit bias in various clinical tasks. Zack et al. (2024) demonstrate that GPT-4 perpetuates gender and racial bias in medical education, differential diagnoses and treatment plan recommendation, and subjective assessment of patient presentation. Yang et al. (2024) show that GPT-3.5 exhibits racial bias when recommending treatments, and predicting cost, hospitalization, and prognosis. Poulain et al. (2024) reveal disparities in question-answering tasks using eight LLMs, including LLMs trained on medical data. Zhang et al. (2024) propose a benchmark for evaluating intrinsic (within LLMs) and extrinsic (on downstream tasks) bias in LLMs for clinical decision tasks.
Motivated by these findings, our work investigates how demographics are encoded by LLMs when they are performing clinical tasks. We have shown that such representations are localized, and that we can effectively control patient demographics by intervening on these representations.
@@
Causal methods have been used to localize demographic information in language models in the general domain. For example, Vig et al. (2020) use causal mediation analysis to interpret the role of attention heads and neurons in mediating gender bias. Chintam et al. (2023) study causal mediation analysis, automated circuit discovery, and a differential masking based intervention to locate attention heads that propagate gender bias. Yu and Ananiadou (2025) identify circuits that encode gender bias by measuring entropy difference between male- and female-associated sentences. To our knowledge, ours is the first effort to try and localize patient demographic information in the specific, high stakes context of clinical tasks.
@@ 6 Conclusions
This work has several important limitations. First, we focussed our analyses on a single open-source LLM, OLMo-7B-Instruct; other LLMs may encode patient demographic information differently. Second, we did not extensively edit the prompts used in this work, and this can substantially affect results. Third, we took a simplistic view of illustrative demographic categories, and in particular—following prior related analyses (Zack et al., 2024)—focussed on patients conforming to binary gender categories; future work might extend this to be more inclusive in analyses.
@@ A.1 Rewrite Score
Layers are numbered from 00 to $L-1$, where $L$ is the number of layers in the LLM.
@@
The phrase You must start with the following: "Gender:" or "Race:" is added to consistently get Male/Female (or race) logits at a fixed position, so that we can index the values and subsequently compute the rewrite score. Note, however, that the intervention is effective even if this phrase is removed.
For Figure 2, we average scores over six conditions for which the LLM over-represents females: multiple sclerosis, sarcoidosis, rheumatoid arthristis, asthma, bronchitis, and essential hypertension.
For certain clinical conditions, such as rheumatoid arthritis (Figure 9(a)) and sarcoidosis (Figure 9(b)), patching at an earlier subtoken of the condition is more effective. This phenomenon is also observed by Meng et al. (2022) perhaps because the model recognizes the condition when it sees “rheumatoid” (or “sar” and “coi”) before seeing “arthritis” or (“osis”) (Feucht et al., 2024).
(a) rheumatoid arthritis
(b) sarcoidosis
Figure 9: Rewrite score distribution for rheumatoid arthritis and sarcoidosis. Patching at an earlier subtoken is more effective.

### v2

@@ Abstract
We know from prior work that LLMs encode social biases, and that this manifests in clinical tasks (Gerszberg, 2024; Zack et al., 2024; Zhang et al., 2020). In this work we adopt tools from mechanistic interpretability to unveil sociodemographic representations and biases within LLMs in the context of healthcare. Specifically, we ask: Can we identify activations within LLMs that encode sociodemographic information (e.g., gender, race)? We find that, in three open weight LLMs, gender information is highly localized in MLP layers and can be reliably manipulated at inference time via patching. Such interventions can surgically alter generated clinical vignettes for specific conditions, and also influence downstream clinical predictions which correlate with gender, e.g., patient risk of depression. We find that representation of patient race is somewhat more distributed, but can also be intervened upon, to a degree. To our knowledge, this is the first application of mechanistic interpretability methods to LLMs for healthcare  [^1] Our code is available at https://github.com/hibaahsan/interp-healthcare-bias/.
@@ 1 Introduction
In this work, we ask: Is the internal LLM encoding of patient demographics like gender and race localized? And, can we intervene upon this? To answer this we perform activation patching (Heimersheim and Nanda, 2024) in the context of clinical vignette generation. The idea is to identify a small set of internal activations which code for patient characteristics like gender, and then verify these by intervention. We offer the following contributions:
(1) We find that gender [^2] Following prior work Zack et al. (2024), we use “gender” instead of “sex” because of the ambiguity in how LLMs use the terms “male” and “female” (biological vs sociocultural). Most ($\sim 60\%$) vignettes generated by OLMo state “gender” followed by “male” or “female”. Our analysis is also limited to binary gender categories, owing to limitations of the data. information is highly localized in MLP layers. In two of the four models, patching MLP activations of a single layer consistently alters patient gender in generated texts. Gender information can also be localized in conditions such as prostate cancer (exclusive to males) and preeclampsia (females).
@@ 2.1 Vignette Generation
Is the encoding of patient gender localized or distributed in the LLM? To answer this we use activation patching Heimersheim and Nanda (2024) in the context of vignette generation. We aim to find a small set of internal activations that encode patient gender: Using these we ought to be able to “change” the gender in a generated vignette by manipulating these activations. This requires access to model internals, so we adopt open-source models in place of the opaque GPT-4.
We focus on the OLMo-7B-Instruct model throughout this paper, but show that our findings generalize across models and scale with three additional models: Llama-3.1-8B-Instruct (Grattafiori et al., 2024), Gemma-2-9B-it (Riviere et al., 2024), and OLMo-2-0325-32B-Instruct (OLMo et al., 2024).
@@
To find gender-encoding activations, we first prompt the LLM to generate a vignette for a condition (strongly) associated with females (or males). Specifically, following prior work Zack et al. (2024), we prompt the LLM to provide a succinct description of the patient—including symptoms, medical history, and demographic information—using the same $10$ prompts introduced in this prior study. One of the prompts, e.g., is:
@@
Figure 2: Rewrite score distribution averaged over six conditions for which OLMo over-represents females. Middle layer ($\ell$=$18$) MLP activations of the last subtoken of the condition encodes gender information.
@@ 2.2 Activation Patching
Where $p(\text{`Male'})$ is the probability of generating the token ‘Male’ for gender when prompting using $x_{\text{vignette}}$ before intervention and $p_{*}(\text{`Male'})$ is the probability of generating ‘Male’ after it. [^5] We append the phrase, You must start with the following: ‘Gender’: to ensure that the next token generated is “Male” or “Female”. Note that the intervention is effective even when the phrase is removed.
@@
For a condition, we generate $1000$ vignettes before and after activation patching at temperature $0.7$. We also experiment with scaling up the MLP activations by a factor, $c$. Figure 3(a) shows the ratio of male vignettes after patching for sarcoidosis and MS. Patching is effective when scaled ($c\geq$22), flipping the gender for all vignettes.
Model Condition Before w/o S w/ S
Llama-3.1-8B-I MS $0.07$ $0.23$ $1.0$
Sarcoidosis $0.06$ $0.19$ $1.0$
Gemma-2-9B-I MS $0.02$ $0.83$ $0.85$
Sarcoidosis $0.07$ $0.92$ $0.92$
OLMo-2-32B-I* MS $0.10$ $0.96$ $0.96$
Table 3: Ratio of male-patient vignettes before and after activation patching. w/o S: without scaling, w/ S: with scaling. *Females were not over-represented for sarcoidosis in OLMo-32B generations.
Table 3 shows results for other models. We observe a similar pattern in Llama-3.1-8B-Instruct: Patching (with scaling) flips the gender to male $100\%$ of the time. Similarly, patching in OLMo-2-32B-Instruct yields vignettes with male patients 96% of the time (scaling is irrelevant here). And in Gemma-2-9B-it, the fraction of males after patching is $0.83$ and $0.92$ for MS and Sarcoidosis, respectively; this is less extreme than other models but a dramatic change nonetheless. [^6] Scaling (up to $c=20$) makes no difference here (see Appendix B for rewrite score plots).
@@
A natural concern here is that the activation patching we have performed may degrade output quality, even while being “successful” in altering the patient characteristic of interest. To assess if this is the case, we compute the perplexity using Llama-3.1-8B [^7] We use Llama as an external judge of perplexity rather than OLMo since the latter generated the text and so likely finds it high likelihood (Panickssery et al., ). Grattafiori et al. (2024) of $500$ vignettes LLM generated before and after patching. Figure 3(b) reports average perplexities as a function of $c$ for sarcoidosis and MS. Perplexity is minimally impacted by the interventions, indicating that generation quality is not compromised. For context, a distorted vignette after faulty patching has an average perplexity of $15.54$ (Appendix 8).
@@ 2.3 Sexed Conditions
We have shown that we can extract a “male activation” which can consistently induce “maleness” via patching in clinical vignettes. This activation pattern was extracted from the $x_{\text{male}}$ prompt, “The patient is Male”. Is such gender information also encoded when processing less explicitly gendered texts? Here we consider the case of conditions which are inherently sexed. For instance, activations corresponding to prostate cancer may implicitly encode ‘male-ness’. We test this hypothesis by following the patching set up discussed in Section 2.2, but change $x_{\text{male}}$ from ‘The patient is Male’ to $x_{\text{vignette}}$ for ‘prostate cancer’. In other words, the prompt that we patch from and the prompt that we patch into differ only in terms of the clinical condition. We patch MLP activations of ‘prostate’ to the last sub-token of the condition in $x_{\text{vignette}}$ at layer $18$. We observe the same phenomenon as shown in Figure 3(a): Patching after scaling activations alters the gender $100\%$ of the time. We also find that the “maleness” patch generalizes to non-clinical domains as well. See Appendix A.3 for details.
@@
Figure 4: Rewrite score distribution for hepatitis B and sarcoidosis. Early (layer 44) as well as middle (layer 182018-20) MLP layers affect racial distribution.
@@ 3 Localizing Race
Model Condition w/o S w/ S SW
Llama-3.1-8B-I* Hepatitis B $0.16$ $0.66$ $0.96$
Gemma-2-9B-I* Sarcoidosis $0.23$ $0.24$ $0.91$
OLMo-2-32B-I Hepatitis B $0.26$ $0.24$ $0.78$
Sarcoidosis $0.06$ $0.08$ $0.76$
Table 4: Proportion of target race after patching. w/o S: without scaling, w/ S: with scaling, SW: sliding window. These are averages over the three target races. *Gemma and Llama did not exhibit skewed racial distributions for hepatitis B and sarcoidosis, respectively.
@@
Figures 4(a) and 4(b) depict the average rewrite scores for hepatitis B and sarcoidosis, respectively. Patching in early ($\ell$=44) as well as middle ($\ell$=182018-20) layers affects racial distribution. Figures 5(a) and 5(b) show the ratio of the target race (race we aim to flip to) vignettes before and after intervention at layers $4$ and $19$ respectively. Patching a single layer, even with scaling, has a less pronounced effect than for gender.
Does patching multiple layers at a time help? We explore sliding window patching Meng et al. (2022), in which neighboring layers are simultaneously patched. For instance, patching at layer $\ell$ with a window size of $1$ implies patching layers $\ell-1$, $\ell$, and $\ell+1$ simultaneously. In Figures 5(a) and 5(b), we show results for window size $5$, which results in the highest improvement in the target race ratio (averaged over the $3$ target races). We report results of window sizes $1$ and $3$ in Tables 14 and 15. Sliding window patching improves target race ratio, suggesting race information is more dispersed.
Table 4 shows results for other models. We observe the same phenomenon in Gemma-2-9B-it and Olmo-2-32B-Instruct: patching a single layer, even with scaling, has a less pronounced effect. Sliding window considerably improves target race ratio. In Llama-3.1-8B-Instruct, patching even a single layer after scaling has a pronounced effect (see Appendix B for details).
As we did for gender, we compute average perplexity under Llama-3.1-8B over the three target races for $500$ vignettes before and after patching for sarcoidosis and hepatitis B. Perplexities before and after patching are comparable (Figure 8).
@@ 4.1 Depression Risk
Similar to the setup in Sections 2 and 3, we extract the ‘male’ and ‘female’ representations using two simple source prompts, “The patient is Male” and “The patient is Female”, respectively. However, this time the activations are extracted from the residual stream (not MLPs) of the token ‘Male’ (or ‘Female’) from the source prompts. Also, the extracted activations are patched at the last token of the target prompt (see Appendix C for examples of the prompt). We choose residual stream since they capture more global information (Geva et al., 2020), allowing us to go beyond a single clinical condition and intervene on notes that typically contain several conditions and confounding variables. Our choice of last token is informed by and aligns with the claims of previous research Marks and Tegmark (2023); Todd et al. (2023) suggesting that the last token representation encodes information about the entire prompt. We patch the target prompt at layer $18$, and scale the activations with a factor of $2$; we picked these values using a set of $100$ BHCs.
@@ Results
Instruction-tuned LLMs are sensitive to instruction phrasings Sun et al. (2023); Ceballos-Arroyo et al. (2024). To ensure our results are robust, we perform the intervention on four different target prompts (see Appendix 13) to elicit risk of depression prediction. Table 6 reports the difference in risk prediction averaged over four prompts for each demographic. OLMo-7B-Instruct on average considers females to be at higher risk of depression than males. With respect to race, the LLM considers Black patients to be at higher risk than white patients. Table 5 shows some sample outputs. While implicit and explicit biases may manifest differently, we observe the same trend in disparity with explicit prompts as well (Appendix C).
@@ 4.2 Differential Diagnosis
We select one case each to study disparities between male/female patients and Black/white patients (see Appendix D for prompts and cases). We follow the set up described in Section 4.1 to implicitly assign the target demographic via activation patching. We use the same source prompts, target layers, and scaling factor.
@@ Results
We sample $500$ differential lists at temperature $0.7$ and check for significant difference in the rank of the correct diagnosis (between male/female and Black/white patients) using Mann-Whitney test. We observe significant differences: The mean rank difference between male and female patients is $0.24$ (p=p=$0.004$), and between Black and white patients is $0.08$ (pp=$0.02$). Figure 7 shows the rank distribution of the correct diagnosis when gender is explicitly stated and when patched; the trend in rank difference is similar whether gender is explicitly or implicitly assigned. See Fig 16 for race.
@@ 5 Related Work
Recent works have shown that LLMs exhibit bias in various clinical tasks. Zack et al. (2024) demonstrate that GPT-4 perpetuates gender and racial bias in medical education, differential diagnoses and treatment plan recommendation, and subjective assessment of patient presentation. Yang et al. (2024) show that GPT-3.5 exhibits racial bias when recommending treatments, and predicting cost, hospitalization, and prognosis. Poulain et al. (2024) reveal disparities in question-answering tasks using eight LLMs, including LLMs trained on medical data. Zhang et al. (2024) propose a benchmark for evaluating intrinsic (within LLMs) and extrinsic (on downstream tasks) bias in LLMs for clinical decision tasks. Zhang et al. (2020) and Kim et al. (2023) quantify biases in domain-adapted masked LMs. Xie et al. (2024) demonstrate that LMs exhibit racial and LGBTQ+ biases using bias benchmarks adapted to the healthcare domain. They further conduct an analysis of debiasing techniques to reduce such biases.
Our work investigates how demographics are encoded by LLMs when they perform clinical tasks – we have shown that such representations are localized. In addition, one can control demographics by intervening on these representations.
@@
Several works have looked at localizing bias in model representations in the general domain. Liang et al. (2020) estimate a bias subspace for sentence representations (generated using a predefined list of bias-sensitive words) using Principal Component Analysis (PCA) (Abdi and Williams, 2010). Ravfogel et al. (2020) train linear probes predictive of the bias attribute to identify a bias subspace. Liang et al. (2021) further extend these works by automatically identifying bias-sensitive words and adopting Ravfogel et al. (2020) for autoregressive generation.
Causal methods have also been used to localize demographic information in language models in the general domain. For example, Vig et al. (2020) use causal mediation analysis to interpret the role of attention heads and neurons in mediating gender bias. Chintam et al. (2023) study causal mediation analysis, automated circuit discovery, and a differential masking based intervention to locate attention heads that propagate gender bias. Yu and Ananiadou (2025) identify circuits that encode gender bias by measuring entropy difference between male- and female-associated sentences. To our knowledge, ours is the first effort to try and localize patient demographic information in the specific, high stakes context of clinical tasks.
@@ 6 Conclusions
This work has several important limitations. First, we did not extensively edit the prompts used in this work, and this can substantially affect results. Second, we took a simplistic view of illustrative demographic categories, and in particular—following prior related analyses (Zack et al., 2024)—focussed on patients conforming to binary gender categories; future work might extend this to be more inclusive in analyses.
7 Acknowledgments
The authors would like to thank Hye Sun Yun and Sheridan Feucht for their feedback on the paper. We acknowledge funding from National Institutes of Health (NIH) under award number R01LM013772 and support from Open Philanthropy.
@@ A.1 Rewrite Score
The phrase You must start with the following: "Gender:" or "Race:" is added to consistently get Male/Female (or race) logits at a fixed position, so that we can index the values and subsequently compute the rewrite score. Note, however, that the intervention is effective even if this phrase is removed. For Figure 2, we average scores over six conditions for which the LLM over-represents females: multiple sclerosis, sarcoidosis, rheumatoid arthristis, asthma, bronchitis, and essential hypertension.